# Large Language Models Miss the Multi-Agent Mark

### v3

@@ 3 Environment Design: MAS LLMs Environments are LLM-centric
As anticipated in the Introduction, MAS traditionally model the environment with few assumptions regarding the architectures of the agents that will populate it. On the other hand, MAS LLMs centre around LLM-powered agents, overlooking the interoperability with agents that do not conform to them. Furthermore, the intrinsic limitations of LLMs, such as their proneness to hallucinations (liu2024exploringevaluatinghallucinationsllmpowered, ; zhou2025exploringtheproblem, ), lack of determinism, and long-term memory persistency (Ouyang_2025, ) may hinder the development of the field, in particular in scenarios where safety and time (as measured in seconds wooldridge ) are paramount.
@@
In MAS, an environment is the external context that agents perceive and react to, and can be characterised across five dimensions (russel2010, ; weyns2004environmentsfor, ). Observability determines the amount of information an agent has about the environment and other agents. Predictability states an environment’s determinism level (from fully deterministic to stochastic). Temporality concerns the dependency between successive environment states. Evolution determines how the environment changes with time (i.e., from discrete or continuous). Manipulation represents the degrees of freedom agents have in interacting and modifying the environment.
@@
Of around $110$ MAS LLMs articles). Furthermore, the vast majority of LLMs articles (see Figure 1, bottom-right) employ textual representations. [^4] As communication systems are not intrinsically part of the environment, we further elaborate on the non-sustainability of text as the preferred medium of communication in large networks of MAS LLMs in Section 4.
@@
In a fully observable setting, agents receive the same, shared representations of the environment; conversely, partially observable environments provide partial, unique representations to each agent. In MAS LLMs observability, whether partial or full, comes with some issues due to the nature of LLMs: in fact, these models notoriously fail at inferring other agents’ beliefs, desires, and intentions (sclar2023mindinglanguagemodelslack, ; shapira2023cleverhansneuraltheory, ; ullman2023largelanguagemodelsfail, ). As discussed in Section 2, LLMs tend to integrate all the information they have access to without distinguishing what they know as opposed to what they know what other agents know, the so called problem of $k^{\text{th}}$ order beliefs (premack1978does, ). In this sense, a centralised MAS LLMs where an LLM orchestrates and monitors the others may incur issues regarding attribute who did what or what an agent wants to achieve. Even systems with two LLMs suffer from these issues (li2023camel, ), as we later discuss in the last paragraph of this section.
@@
While most MAS LLMs papers assume the environment is deterministic and can be modified (i.e., it changes according to the actions the LLMs make), the intrinsic non-determinism of LLMs flaws this setting. In terms of safety, one can design specific procedures, such as safety mechanisms and guardrail measures, that deterministically trigger when particular events happen; on the other hand, non-deterministic LLMs provide no guarantees they will behave accordingly (chen2024agentpoisonredteamingllmagents, ). Even setting the temperature of an LLM to zero or sampling deterministically their strategies is not a solution, as many LLMs are known to behave always non-deterministically (lamalfa2023jair, ; Ouyang_2025, ).
@@
In two popular frameworks, Camel (li2023camel, ) and MetaAgent (li2023metaagentssimulatinginteractionshuman, ), the authors illustrate how their models hallucinate and underperform when MAS LLMs are specialised to perform a subtask. In Camel, two LLMs autonomously generate prompts to solve complex tasks, reducing reliance on user input. The authors notice that LLMs can inadvertently swap their roles, generate repeated or non-useful instructions or get stuck in infinite message exchanges. MetaAgent focuses on collaborative agents that accomplish coordination tasks. Through perception, memory, reasoning, and execution modules, LLMs interact with the environment, store valuable information, and learn rewarding skills. The authors show that LLMs deviate from their predefined identities, hallucinating their competence and thus compromising collaboration.
@@
We argue that these problems are a byproduct of the overreliance on natural language and free-text as the reference medium for coordinating agents. Natural language is inherently ambiguous and prone to misinterpretation (battle2024unreasonableeffectivenesseccentricautomatic, ; li2025alwaysniceconfidentwrong, ), suffers from information loss (Li_2023, ), and requires high costs for storing and retrieving information. In the following paragraph, we propose some research directions to mitigate these issues.
@@
Research directions. We propose the following research directions to address the issues with textual environments. First, we argue it is crucial to explore multi-modal environments where the stimuli are turned into actionable steps without intermediate translation into natural language (Jeong_2025, ; liu2024lmagentlargescalemultimodalagents, ; zhang2024speechagentshumancommunicationsimulationmultimodal, ). The intuition is that the fewer mediations a stimulus goes through, the less likely the signal will contain distortions or introduce errors.
@@
Furthermore, using structured formats (taking inspiration from performative agentic languages such as KQML (kqml1994, )) can remove the ambiguities inherent in natural language communication. Last but not least, integrating LLMs and formal planners or neuro-symbolic methods can provide guarantees that precise actions will be carried out correctly: LLMs excel at extracting a task’s specifics from noisy and incomplete specifications, while formal methods can provide plans that are guaranteed to achieve the goal.
@@ 4 Coordination and Communication Issues in MAS LLMs
Coordination and the lack of asynchronous MAS LLMs. As described in Section 1, asynchronicity is a key component of multi-agent concurrent systems: in its basic form, concurrent systems encompass multiple, diverse tasks executed simultaneously, without assuming when they start or end. Examples of notorious problems that can only be modelled with asynchronicity are those that necessitate concurrent algorithms to be solved (10.5555/2029110, ; diaz1981formalization, ): these include classic problems such as the “dining philosophers”, handshaking protocols, etc. Asynchronicity also arises in many practical scenarios, such as email and chat exchanges and database management access. Many real-world scenarios cannot be modelled without asynchronicity and require simplification (e.g., by assuming agents act sequentially through an orchestrator).
Notably, while most closed-source LLM providers offer asynchronous APIs,.
@@
For example, we consider the case of AutoGen, an influential MAS LLMs framework (wu2023autogen, ) that enables building LLM applications through multiple interacting agents. Although AutoGen supports asynchronous calls, [^6] https://github.com/microsoft/autogen?tab=readme-ov-file#web-browsing-agent-team developers must define asynchronous calls for each action and event; asynchronous programming with synchronous languages is well-known to be prone to bugs that impede the system from being fully asynchronous. As we expand later, we argue for a reverse approach. Every agent and environment should be natively asynchronous to be considered as a MAS, [^7] This core characteristic is discussed in detail in (wooldridge, ), Chapter $1.3$. with sequential calls being the exception, rather than the norm.
Communications methods in MAS LLMs. Traditionally, there are three levels at which (Multi-Agent) communication is analysed (austin1962how, ). An utterance that conveys some meaning may have to the hearer no intended effect (an illocutionary act, e.g., "the sky is blue"), some meaning intended to warn (a perlocutionary act, "a train is passing"), or some meaning that acts as a request (a performative, e.g., "please open the window"). These distinctions, alongside the so-called rules of conversation (implicature, the Gricean maxims (Grice1975-GRILAC-6, ), etc.), constitute the foundation of agents’ communication in MAS (berna2004communication, ; jang2005efficient, ; saunders1996evolution, ).
In MAS, illocutionary conversations are usually handled using descriptive and structured languages, such as JSON and RDF. Their purpose is to exchange information, not to perform or request actions. On the other hand, KQML (kqml1994, ), FIPA’s ACL, [^8] http://www.fipa.org/specs/fipa00061/XC00061D.html and rational programming and Agent-Oriented Programming (shoham1993agent, ) are examples of consolidated performative standards that describe agents’ beliefs, commitments, and actions. For instance, for an agent that exchanges a file with another and asks to summarise it, the MAS literature proposes using a structured language for the exchange and a performative query for the summarisation.
When it comes to LLMs, humans interact with them in free-text form; by extension, most MAS LLMs systems adopt natural language as the primary communication medium. While natural language captures the nuances of the principles mentioned above, it comes at the cost of complexity and ambiguity: an utterance such as "a car is coming" may range from being illocutionary (conveying information) to performative (conveying an implicit warning or an order). A few recent works in MAS LLMs propose to handle handshakes and errors with natural language communications; any other communication that routines and protocols can implement should otherwise favour structured languages (marro2024scalablecommunicationprotocolnetworks, ).
@@
Research directions. In terms of asynchronicity, we devise two complementary research directions that are worth investigating: on the one hand, frameworks that model MAS as asynchronous systems (CORNFORTH200570, ) should be adapted to MAS LLMs; the capacity of modelling asynchronous systems can provide insights and guarantees into their long-term evolution. One example is that described in (adobbati2025asynchronousmultiagentsystemspetri, ), which models MAS with Petri nets, a model of computation that is inherently asynchronous, enabling analyses on reachability, boundedness and invariance of the system.
On the other hand, as mentioned above, providers often expose asynchronous LLM APIs, for which we should develop suitable environments. Critical points in this sense are how deadlocks and starvation are handled to ensure a consistent evolution of the environment (ginart2024asynchronoustoolusagerealtime, ; sami2025nexuslightweightscalablemultiagent, ). To summarise our position, we argue for MAS LLM frameworks that are natively asynchronous and where sequential actions are the exception. As regards communication systems, we argue for standard, open-source frameworks to reason and build the key components of any LLMs communication (e.g., aspects such as security, identity preservation, trust, message exchange, etc.).
In line with some recent initiatives, [^9] https://github.com/google/A2A, https://nanda.media.mit.edu/, https://las-wg.org/ we believe the MAS LLMs community should work towards standard agent protocols guided, where needed, by MAS principles. In the spirit of Agent Oriented Programming (weyns2010architecture, ), researchers would have abstract templates that already implement security routines (e.g., a communication between web agents would only happen over a secure channel such as HTTPS).
@@ 5 MAS LLMs do not Quantify the Emergence of Complex Behaviours
In many disciplines, such as system theory, economics, MAS, etc., an emergent behaviour is observed when a complex entity has properties or behaviours that its parts do not have on their own, and emerge only when they interact in a wider whole (altmann2024emergencemultiagentsystemssafety, ; fromm2005types, ). In the context of LLMs, emergence describes the increasing capabilities of a model at varying model size (brown2020languagemodelsfewshotlearners, ; kaplan2020scalinglawsneurallanguage, ). With MAS LLMs, behaviours that cannot be predicted via a static analysis of agents and their environment are also considered emergent Li_2023 ; park2023generative .
@@
Another recent work studies how LLMs build agent societies within a Minecraft environment (al2024projectsidmanyagentsimulations, ). While the work claims that agents can achieve significant milestones towards AI civilisations, [^10] A video showcasing their emergent behaviours, https://www.youtube.com/watch?v=9piFiQJ-mnU. results are primarily observational, i.e., the system is let evolve for a long time and then researchers make their observations on interesting behaviours the LLMs adopt. Other works approach emergence from a similar perspective (park2023generative, ; xie2024largelanguagemodelagents, ), and a body of research studies emergence before the advent of ChatGPT (graesser2020emergentlinguisticphenomenamultiagent, ; lazaridou2020emergentmultiagentcommunicationdeep, ).
@@
In conclusion, in this body of research, we observe (i) no systematic definition of what constitutes an emergent behaviour, without reference to the system being analysed, therefore (ii) a lack of proper benchmarks to quantify them. In contrast, traditional MAS research generally insists on rigorously defined environments, quantifiable objectives, and formal verification methods. Emergence alone, especially when derived from loosely constrained prompts and undefined interaction dynamics, does not suffice to claim long-horizon MAS-level coordination or planning capabilities (turner2022formalizingproblemeffectregularization, ).
Research directions. We argue for a proper definition of emergence and emergent behaviours in open-ended MAS LLMs, where the concept is well established and scientifically falsifiable (altmann2024emergencemultiagentsystemssafety, ; fromm2005types, ).
For example, in economics, emergence is sometimes characterised in its core aspects, and encompasses behaviours that produce outcomes the theory can explain (kuperberg2007two, ): in other words, emergent phenomena are economics phenomena. In this sense, since MAS LLMs share with economics the interest in agents’ behaviour, a characterisation of emergence can encompass phenomena that relate to the agents’ objective functions (e.g., in a system where agents have to maximise productivity, one finds a way to hack the reward function).
@@
While this definition has the advantage of being measurable, it fits the notions of weak and strong emergence in computer science (chalmers2006strong, ) (a weak emergent phenomenon can be derived from the underlying system; a strong phenomenon requires new laws or assumptions), thus making the adaptation to MAS LLMs more straightforward.

### v4

@@ 3 Environment Design: MAS LLMs Environments are LLM-centric
As anticipated in the Introduction, MAS traditionally model the environment with few assumptions regarding the architectures of the agents that will populate it. On the other hand, MAS LLMs centre around LLM-powered agents, overlooking the interoperability with agents that do not conform to them. Furthermore, the intrinsic limitations of LLMs, such as their proneness to hallucinations (liu2024exploringevaluatinghallucinationsllmpowered; zhou2025exploringtheproblem), lack of determinism, and long-term memory persistency (Ouyang_2025) may hinder the development of the field, in particular in scenarios where safety and time (as measured in seconds wooldridge) are paramount.
@@
In MAS, an environment is the external context that agents perceive and react to, and can be characterised across five dimensions (russel2010; weyns2004environmentsfor). Observability determines the amount of information an agent has about the environment and other agents. Predictability states an environment’s determinism level (from fully deterministic to stochastic). Temporality concerns the dependency between successive environment states. Evolution determines how the environment changes with time (i.e., from discrete or continuous). Manipulation represents the degrees of freedom agents have in interacting and modifying the environment.
@@
Of around $110$ MAS LLMs articles published between $2023$ and $2025$ and whose results we summarise in Figure 1, most MAS LLMs operate in partially observable, deterministic, temporally dynamic, discrete environments (li2024surveyonllm; minaee2025largelanguagemodelssurvey) (Further details on the papers in Appendix A). Furthermore, the vast majority of LLMs articles (see Figure 1, bottom-right) employ textual representations. [^4] As communication systems are not intrinsically part of the environment, we further elaborate on the non-sustainability of text as the preferred medium of communication in large networks of MAS LLMs in Section 4.
@@
In a fully observable setting, agents receive the same, shared representations of the environment; conversely, partially observable environments provide partial, unique representations to each agent. In MAS LLMs observability, whether partial or full, comes with some issues due to the nature of LLMs: in fact, these models notoriously fail at inferring other agents’ beliefs, desires, and intentions (sclar2023mindinglanguagemodelslack; shapira2023cleverhansneuraltheory; ullman2023largelanguagemodelsfail). As discussed in Section 2, LLMs tend to integrate all the information they have access to without distinguishing what they know as opposed to what they know what other agents know, the so called problem of $k^{\text{th}}$ order beliefs (premack1978does). In this sense, a centralised MAS LLMs where an LLM orchestrates and monitors the others may incur issues regarding attribute who did what or what an agent wants to achieve. Even systems with two LLMs suffer from these issues (li2023camel), as we later discuss in the last paragraph of this section.
@@
While most MAS LLMs papers assume the environment is deterministic and can be modified (i.e., it changes according to the actions the LLMs make), the intrinsic non-determinism of LLMs flaws this setting. In terms of safety, one can design specific procedures, such as safety mechanisms and guardrail measures, that deterministically trigger when particular events happen; on the other hand, non-deterministic LLMs provide no guarantees they will behave accordingly (chen2024agentpoisonredteamingllmagents). Even setting the temperature of an LLM to zero or sampling deterministically their strategies is not a solution, as many LLMs are known to behave always non-deterministically (lamalfa2023jair; Ouyang_2025).
@@
In two popular frameworks, Camel (li2023camel) and MetaAgent (li2023metaagentssimulatinginteractionshuman), the authors illustrate how their models hallucinate and underperform when MAS LLMs are specialised to perform a subtask. In Camel, two LLMs autonomously generate prompts to solve complex tasks, reducing reliance on user input. The authors notice that LLMs can inadvertently swap their roles, generate repeated or non-useful instructions or get stuck in infinite message exchanges. MetaAgent focuses on collaborative agents that accomplish coordination tasks. Through perception, memory, reasoning, and execution modules, LLMs interact with the environment, store valuable information, and learn rewarding skills. The authors show that LLMs deviate from their predefined identities, hallucinating their competence and thus compromising collaboration.
@@
We argue that these problems are a byproduct of the overreliance on natural language and free-text as the reference medium for coordinating agents. Natural language is inherently ambiguous and prone to misinterpretation (battle2024unreasonableeffectivenesseccentricautomatic; li2025alwaysniceconfidentwrong), suffers from information loss (Li_2023), and requires high costs for storing and retrieving information. In the following paragraph, we propose some research directions to mitigate these issues.
@@
Research directions. We propose the following research directions to address the issues with textual environments. First, we argue it is crucial to explore multi-modal environments where the stimuli are turned into actionable steps without intermediate translation into natural language (Jeong_2025; liu2024lmagentlargescalemultimodalagents; zhang2024speechagentshumancommunicationsimulationmultimodal). The intuition is that the fewer mediations a stimulus goes through, the less likely the signal will contain distortions or introduce errors.
@@
Furthermore, using structured formats (taking inspiration from performative agentic languages such as KQML (kqml1994)) can remove the ambiguities inherent in natural language communication. Last but not least, integrating LLMs and formal planners or neuro-symbolic methods can provide guarantees that precise actions will be carried out correctly: LLMs excel at extracting a task’s specifics from noisy and incomplete specifications, while formal methods can provide plans that are guaranteed to achieve the goal.
@@ 4 Coordination and Communication Issues in MAS LLMs
Coordination and the lack of asynchronous MAS LLMs. As described in Section 1, asynchronicity is a key component of multi-agent concurrent systems: in its basic form, concurrent systems encompass multiple, diverse tasks executed simultaneously, without assuming when they start or end. Examples of notorious problems that can only be modelled with asynchronicity are those that necessitate concurrent algorithms to be solved (10.5555/2029110; diaz1981formalization): these include classic problems such as the “dining philosophers”, handshaking protocols, etc. Asynchronicity also arises in many practical scenarios, such as email and chat exchanges and database management access. Many real-world scenarios cannot be modelled without asynchronicity and require simplification (e.g., by assuming agents act sequentially through an orchestrator).
Notably, while most closed-source LLM providers offer asynchronous APIs, agents employing LLMs tend to be predominantly used in synchronous or parallel fashion (li2024agentsneed).
@@
For example, we consider the case of AutoGen, an influential MAS LLMs framework (wu2023autogen) that enables building LLM applications through multiple interacting agents. Although AutoGen supports asynchronous calls, [^6] https://github.com/microsoft/autogen?tab=readme-ov-file#web-browsing-agent-team developers must define asynchronous calls for each action and event; asynchronous programming with synchronous languages is well-known to be prone to bugs that impede the system from being fully asynchronous. As we expand later, we argue for a reverse approach. Every agent and environment should be natively asynchronous to be considered as a MAS, [^7] This core characteristic is discussed in detail in (wooldridge), Chapter $1.3$. with sequential calls being the exception, rather than the norm.
Communications methods in MAS LLMs. Traditionally, there are three levels at which (Multi-Agent) communication is analysed (austin1962how). An utterance that conveys some meaning may have to the hearer no intended effect (an illocutionary act, e.g., "the sky is blue"), some meaning intended to warn (a perlocutionary act, "a train is passing"), or some meaning that acts as a request (a performative, e.g., "please open the window"). These distinctions, alongside the so-called rules of conversation (implicature, the Gricean maxims (Grice1975-GRILAC-6), etc.), constitute the foundation of agents’ communication in MAS (berna2004communication; jang2005efficient; saunders1996evolution).
In MAS, illocutionary conversations are usually handled using descriptive and structured languages, such as JSON and RDF. Their purpose is to exchange information, not to perform or request actions. On the other hand, KQML (kqml1994), FIPA’s ACL, [^8] http://www.fipa.org/specs/fipa00061/XC00061D.html and rational programming and Agent-Oriented Programming (shoham1993agent) are examples of consolidated performative standards that describe agents’ beliefs, commitments, and actions. For instance, for an agent that exchanges a file with another and asks to summarise it, the MAS literature proposes using a structured language for the exchange and a performative query for the summarisation.
When it comes to LLMs, humans interact with them in free-text form; by extension, most MAS LLMs systems adopt natural language as the primary communication medium. While natural language captures the nuances of the principles mentioned above, it comes at the cost of complexity and ambiguity: an utterance such as "a car is coming" may range from being illocutionary (conveying information) to performative (conveying an implicit warning or an order). A few recent works in MAS LLMs propose to handle handshakes and errors with natural language communications; any other communication that routines and protocols can implement should otherwise favour structured languages (marro2024scalablecommunicationprotocolnetworks).
@@
Research directions. In terms of asynchronicity, we devise two complementary research directions that are worth investigating: on the one hand, frameworks that model MAS as asynchronous systems (CORNFORTH200570) should be adapted to MAS LLMs; the capacity of modelling asynchronous systems can provide insights and guarantees into their long-term evolution. One example is that described in (adobbati2025asynchronousmultiagentsystemspetri), which models MAS with Petri nets, a model of computation that is inherently asynchronous, enabling analyses on reachability, boundedness and invariance of the system.
On the other hand, as mentioned above, providers often expose asynchronous LLM APIs, for which we should develop suitable environments. Critical points in this sense are how deadlocks and starvation are handled to ensure a consistent evolution of the environment (ginart2024asynchronoustoolusagerealtime; sami2025nexuslightweightscalablemultiagent). To summarise our position, we argue for MAS LLM frameworks that are natively asynchronous and where sequential actions are the exception. As regards communication systems, we argue for standard, open-source frameworks to reason and build the key components of any LLMs communication (e.g., aspects such as security, identity preservation, trust, message exchange, etc.).
In line with some recent initiatives, [^9] https://github.com/google/A2A, https://nanda.media.mit.edu/, https://las-wg.org/ we believe the MAS LLMs community should work towards standard agent protocols guided, where needed, by MAS principles. In the spirit of Agent Oriented Programming (weyns2010architecture), researchers would have abstract templates that already implement security routines (e.g., a communication between web agents would only happen over a secure channel such as HTTPS).
@@ 5 MAS LLMs do not Quantify the Emergence of Complex Behaviours
In many disciplines, such as system theory, economics, MAS, etc., an emergent behaviour is observed when a complex entity has properties or behaviours that its parts do not have on their own, and emerge only when they interact in a wider whole (altmann2024emergencemultiagentsystemssafety; fromm2005types). In the context of LLMs, emergence describes the increasing capabilities of a model at varying model size (brown2020languagemodelsfewshotlearners; kaplan2020scalinglawsneurallanguage). With MAS LLMs, behaviours that cannot be predicted via a static analysis of agents and their environment are also considered emergent Li_2023; park2023generative.
@@
Another recent work studies how LLMs build agent societies within a Minecraft environment (al2024projectsidmanyagentsimulations). While the work claims that agents can achieve significant milestones towards AI civilisations, [^10] A video showcasing their emergent behaviours, https://www.youtube.com/watch?v=9piFiQJ-mnU. results are primarily observational, i.e., the system is let evolve for a long time and then researchers make their observations on interesting behaviours the LLMs adopt. Other works approach emergence from a similar perspective (park2023generative; xie2024largelanguagemodelagents), and a body of research studies emergence before the advent of ChatGPT (graesser2020emergentlinguisticphenomenamultiagent; lazaridou2020emergentmultiagentcommunicationdeep).
@@
In conclusion, in this body of research, we observe (i) no systematic definition of what constitutes an emergent behaviour, without reference to the system being analysed, therefore (ii) a lack of proper benchmarks to quantify them. In contrast, traditional MAS research generally insists on rigorously defined environments, quantifiable objectives, and formal verification methods. Emergence alone, especially when derived from loosely constrained prompts and undefined interaction dynamics, does not suffice to claim long-horizon MAS-level coordination or planning capabilities (turner2022formalizingproblemeffectregularization).
Research directions. We argue for a proper definition of emergence and emergent behaviours in open-ended MAS LLMs, where the concept is well established and scientifically falsifiable (altmann2024emergencemultiagentsystemssafety; fromm2005types).
For example, in economics, emergence is sometimes characterised in its core aspects, and encompasses behaviours that produce outcomes the theory can explain (kuperberg2007two): in other words, emergent phenomena are economics phenomena. In this sense, since MAS LLMs share with economics the interest in agents’ behaviour, a characterisation of emergence can encompass phenomena that relate to the agents’ objective functions (e.g., in a system where agents have to maximise productivity, one finds a way to hack the reward function).
@@
While this definition has the advantage of being measurable, it fits the notions of weak and strong emergence in computer science (chalmers2006strong) (a weak emergent phenomenon can be derived from the underlying system; a strong phenomenon requires new laws or assumptions), thus making the adaptation to MAS LLMs more straightforward.